OpenStudy (anonymous):

A bus travels 2800m south along a straight path with an avg. velocity of 24m/s to the south. The bus stops for 24 min., then it travels2100 m south w/average velocity of 20.8 m/s to the south. How long does the total trip last? What is the average velocity for the total trip?

OpenStudy (anonymous):

I have two answers the first one 16 hrs and 26 minutes and the second one172.19 m/s but i don't know if it is right, please help :(

OpenStudy (shane_b):

As with any physics problem, once you've calculated the answer you should always ask yourself: does my answer make sense? It shouldn't make sense that it would take *hours* to go 2800m if you're moving at an average velocity of over 20 m/s...even with a 24 minute delay. Ok, let's work through it. Start by seeing how long the bus to do each leg: \[t_1=\frac{2800m}{24m/s}=116.67s\]\[t_2=\frac{2100m}{20.8m/s}=100.96s\]\[t_3=24min==>24min*60s/min=1440s\] How long does the total trip last?\[t_{total}=t_1+t_2+t_3=116.67s+100.96s+1440s=1657.63s\] What is the average velocity for the total trip? The average velocity will be the total displacement divided by the time it took to make the trip:\[\frac{2800m+2100m}{1657.63s}=2.96m/s\]That should sound about right...since the bus spent most of the trip at the bus stop. <Note that you'll have to adjust the calcs to account for significant figures>